A. Solve the following problems:

1. On the assumption that variables are continuous, compute for the median of the following data:

a. 1, 4, 5, 9, 11

b. 4, 8, 16, 27, 52

c. 7, 26, 31, 36, 46, 65

2. Calculate the range for the following sets of measurements:

a. 3, 6, 7, 10, 2

b. 11, 27, 32, 45, 54

3. Calculate the variance and standard deviation of the following: 13, 18, 21, 19, 16, 15.



1
Expert's answer
2021-12-07T11:31:07-0500

1.

a.


"median=5"


b.


"median=16"

c.


"median=\dfrac{31+36}{2}=33.5"

2.

a.

"Range=10-2=8"

b.


"Range=54-11=43"

3.


"mean=E(X)=\dfrac{1}{6}(13+18+21+19+16+15)"

"=\dfrac{102}{6}=17"

"Var(X)=\sigma^2=\dfrac{1}{6}((13-17)^2+(18-17)^2"

"+(21-17)^2+(19-17)^2+(16-17)^2"

"+(15-17)^2)=\dfrac{42}{6}=7"

"\sigma=\sqrt{\sigma^2}=\sqrt{7}"